# Extracting energy from plunging region of a Kerr-Taub-NUT black hole by magnetic reconnection

### v1

@@ 3 Energy extraction by magnetic reconnections in the Kerr-Taub-NUT spacetime
In Fig. 1 we present effects of black hole spin $a$ and gravitomagnetic charge $l$ on the magnetic field angle. In the Kerr-Taub-NUT spacetime, $\xi$ for the plunging plasmoid is also negative and its absolute value increases as the plasmoid gets closer to the black hole’s horizon, which is similar to those of the Kerr case. The presence of gravitomagnetic charge $l$ decreases the change rate of the magnetic field angle $\xi$ with respect to $r$. The absolute value of $\xi$ at the horizon decreases with the gravitomagnetic charge $l$. In the case $l=0.2$, the magnetic field angle at the horizon approaches a minimum value of $-54^{\circ}$ as $a\rightarrow 1.019$, whose absolute value is less than that in the Kerr case where it tends to $-60^{\circ}$ as $a\rightarrow 1$. These results show that effects of the gravitomagnetic charge on $\xi$ are opposite to those of the black hole spin.
@@
Moreover, the radius of the reconnection location increases with the gravitomagnetic charge $l$ for fixed $\beta^{\phi}$, which is shown in Fig. 3.
@@
Fig. 5 displays the maximum value of $\eta$ under various spins, gravitomagnetic charges and magnetization parameters. We find that the range of $a$ at which the energy extraction efficiency exceeds $1$ decreases with the gravitomagnetic parameter $l$. In other words, in order to attain an equivalent peak efficiency, the Kerr–Taub-NUT black hole demands a spin significantly higher than its Kerr counterpart. Moreover, the maximum value $\eta_{max}$ increases with the magnetization parameter $\sigma$ and almost reaches a constant once $\sigma$ exceeds approximately $100$. This means that one can use $\sigma=100$ to represent cases of extremely high magnetization as in the Kerr case [32]. In the bottom right panel of Fig. 5, we plot the relationship curves between spin and gravitomagnetic charge under the condition that the maximum value of $\eta=1$. In the parameter plane $l-a$, the effective energy extraction (that is, $\eta>1$) occurs only in the region between the curve corresponding to $\eta_{max}=1$ and the one-horizon curve (that is, $a^{2}-l^{2}=M^{2}$). It is notable that for low magnetization $\sigma\simeq 1$, there is nearly no allowed parameter region because the curve $\eta_{max}=1$ overlaps with the boundary without black hole existence.
@@
Figure 6: Changes of the critical radius with black hole spin $a$, magnetization parameter $\sigma$ and NUT charge $l$. The region in red and purple means extracting energy can escape to the infinity and extraction efficiency $\eta$ over $1$. Only between the orange and blue dashed curves can $\eta$ be above 1. The green dashed curve stands for the critical radius for the escaping condition, above which the plasmoid with $\epsilon_{+}$ can escape to the infinity.
@@ 4 Explanation to jet power of source GRS1915+105
Figure 9: The allowed regions in the $l-a$ plane for the Kerr-NUT black hole. The red colored region is obtained from the observed jet power of GRS1915+105 by the magnetic reconnection mechanism with the Lorentz factor $\Gamma=2$ and magnetization $\sigma=40$. The blue region is obtained by the observed radiative efficiency for GRS1915+105 and the blue solid curve corresponds to the case with radiative efficiency $\eta=0.234$. The red solid line is for the central value of the jet power $P_{jet}=42$ while the red dashed lines correspond to error of $0.3dex$.
For the source GRS1915+105, the total power is observed as $P_{jet}=42$ with the error of 0.3 $dex$ for $\Gamma=2$ and radiative efficiency is in the range $\eta>0.234$ from the data listed in [50]. With these data, in Fig. 9, we present the allowed region in the parameter plane $l-a$. We find that for $\sigma=40$ the allowed parameter region originating from the jet power has an intersection with the region from the radiative efficiency. This implies that through energy extraction via magnetic reconnection mechanism the Kerr-Taub-NUT metric is found to simultaneously explain the observed jet power and radiative efficiency of GRS 1915+105, which is different from that obtained by the Blandford-Znajek mechanism.
@@ 5 Summary
We have studied the energy extraction from a Kerr-Taub-NUT black hole via magnetic reconnections occurring in the plunging region. For plunging plasmoids, the presence of gravitomagnetic charge $l$ decreases the absolute value of magnetic field angle $\xi$ at the horizon and its change rate with respect to $r$, which means that effects of the gravitomagnetic charge $l$ on the magnetic field angle are opposite to those of the black hole spin. Moreover, the gravitomagnetic charge $l$ is not conducive to the generation of negative energy $\epsilon_{-}$ and reduces the efficiency of energy extraction, while increasing the radial coordinate at which the minimum value of $\epsilon_{-}$ appears. Furthermore, the gravitomagnetic parameter $l$ shortens and decreases the range of values for the black hole spin $a$ in which the energy extraction efficiency $\eta>1$, which means that to attain an equivalent peak efficiency, the Kerr–Taub-NUT black hole demands a spin significantly higher than its Kerr counterpart. We also investigate the critical radius $r_{c}$ obtained by the escape condition of the plasmoid with $\epsilon_{+}$ and found that it increases with the NUT charge $l$. Overall, the gravitomagnetic charge suppresses the energy extraction process through magnetic reconnection, which is opposite to those of the black hole spin and the magnetization parameter.

### v2

@@ 3 Energy extraction by magnetic reconnections in the Kerr-Taub-NUT spacetime
In Fig. 1 we present effects of the black hole spin $a$ and gravitomagnetic charge $l$ on the magnetic field angle. In the Kerr-Taub-NUT spacetime, the angle $\xi$ for the plunging plasmoid is also negative and its absolute value increases as the plasmoid gets closer to the black hole’s horizon, which is similar to those of the Kerr case. The presence of gravitomagnetic charge $l$ decreases the change rate of the magnetic field angle $\xi$ with respect to $r$. The absolute value of $\xi$ at the horizon decreases with the gravitomagnetic charge $l$. In the case $l=0.2$, the magnetic field angle at the horizon approaches a minimum value of $54^{\circ}$ as $a\rightarrow 1.019$, whose absolute value is less than that in the Kerr case where it tends to $60^{\circ}$ as $a\rightarrow 1$. These results show that effects of the gravitomagnetic charge on the angle $\xi$ are opposite to those of the black hole spin.
@@
Fig. 5 displays the maximum value of $\eta$ under various spins, gravitomagnetic charges and magnetization parameters. We find that the range of $a$ at which the energy extraction efficiency exceeds $1$ decreases with the gravitomagnetic parameter $l$. In other words, in order to attain an equivalent peak efficiency, the Kerr–Taub-NUT black hole demands a spin significantly higher than its Kerr counterpart. Moreover, the maximum value $\eta_{max}$ increases with the magnetization parameter $\sigma$ and almost reaches a constant once $\sigma$ exceeds approximately $100$. This means that one can use $\sigma=100$ to represent cases of extremely high magnetization as in the Kerr case [63]. In the bottom right panel of Fig. 5, we plot the relationship curves between spin and gravitomagnetic charge under the condition that the maximum value of $\eta=1$. In the parameter plane lal-a, the effective energy extraction (that is, $\eta>1$) occurs only in the region between the curve corresponding to $\eta_{max}=1$ and the one-horizon curve (that is, $a^{2}-l^{2}=M^{2}$). It is notable that for low magnetization $\sigma\simeq 1$, there is nearly no allowed parameter region because the curve $\eta_{max}=1$ overlaps with the boundary without black hole existence.
@@
Figure 6: Changes of the critical radius with black hole spin $a$, magnetization parameter $\sigma$ and NUT charge $l$. The region in red and purple means extracting energy can escape to the infinity and extraction efficiency $\eta$ over 11. Only between the orange and blue dashed curves can $\eta$ be above 1. The green dashed curve stands for the critical radius for the escaping condition, above which the plasmoid with $\epsilon_{+}$ can escape to the infinity.
@@ 4 Explanation to jet power of source GRS1915+105
Figure 9: The allowed regions in the lal-a plane for the Kerr-NUT black hole. The red colored region is obtained from the observed jet power of GRS1915+105 by the magnetic reconnection mechanism with the Lorentz factor $\Gamma=2$ and magnetization $\sigma=40$. The blue region is obtained by the observed radiative efficiency for GRS1915+105 and the blue solid curve corresponds to the case with radiative efficiency $\eta=0.234$. The red solid line is for the central value of the jet power $P_{jet}=42$ while the red dashed lines correspond to error of $0.3dex$.
For the source GRS1915+105, the total power is observed as $P_{jet}=42$ with the error of 0.3 dexdex for $\Gamma=2$ and radiative efficiency is in the range $\eta>0.234$ from the data listed in [81]. With these data, in Fig. 9, we present the allowed region in the parameter plane lal-a. We find that for $\sigma=40$ the allowed parameter region originating from the jet power has an intersection with the region from the radiative efficiency. This implies that through energy extraction via magnetic reconnection mechanism the Kerr-Taub-NUT metric is found to simultaneously explain the observed jet power and radiative efficiency of GRS 1915+105, which is different from that obtained by the Blandford-Znajek mechanism.
@@ 5 Summary
We have studied the energy extraction from a Kerr-Taub-NUT black hole via magnetic reconnections occurring in the plunging region. For plunging plasmoids, the presence of gravitomagnetic charge $l$ decreases the absolute value of magnetic field angle $\xi$ at the horizon and its change rate with respect to $r$, which means that effects of the gravitomagnetic charge $l$ on the magnetic field angle are opposite to those of the black hole spin. Moreover, the gravitomagnetic charge $l$ is not conducive to the generation of negative energy $\epsilon_{-}$ and reduces the efficiency of energy extraction, while increasing the radial coordinate at which the minimum value of $\epsilon_{-}$ appears. Furthermore, the gravitomagnetic parameter $l$ shortens and decreases the range of values for the black hole spin $a$ in which the energy extraction efficiency $\eta>1$, which means that to attain an equivalent peak efficiency, the Kerr–Taub-NUT black hole demands a spin significantly higher than its Kerr counterpart. We also investigated the critical radius $r_{c}$ obtained by the escape condition of the plasmoid with $\epsilon_{+}$ and found that it increases with the NUT charge $l$. Overall, the gravitomagnetic charge suppresses the energy extraction process through magnetic reconnection, which is opposite to those of the black hole spin and the magnetization parameter.